Answer the following in mm units

a) The voltage across a parallel plate capacitor that has a plate separation equal to 0.780 mm is 1.10 KV. The capacitor is disconnected from the voltage source and the separation between the plates is increased until the energy stored in the capacitor has been doubled. Determine the final separation between the plates.
b) If a conducting sphere is to be charged to a potential of 5700 V, what is the smallest possible radius of the sphere such that the electric field will not exceed the dielectric strength of air?
Answer #1

b) Here, V= 5700 volt =(Q/4)×(1/r)

E=3000 volt/mm =(Q/4)×(1/r^2)

we get from the two equation,

(5700/3000)=(1/r)/(1/r^2)

=> (57/30)= r

=> r= 1.9 mm

hope your questions have been answered proprerly.

happy HomeworkLibing.